OpenStudy (anonymous):

cos(37.5)sin(7.5) cos(105)-cos(15) #s are in degrees

OpenStudy (anonymous):

cos(30+7.5)= cos(30)cos(7.5) - Sin(30)Sin(7.5)

hero (hero):

cos(105)-cos(15) = -sqrt(6)/2